# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ 6.2. Post-action Stage
The post-action stage closes the loop after an action is executed. Given execution feedback, the agent evaluates whether the outcome satisfies user-conditioned objectives, applies corrective adjustments when mismatches arise (execution-time correction), and produces final outputs that conform to user preferences and quality criteria (outcome realization).
@@ 6.2.1. Execution-Time Correction
A failed execution commonly produces feedback signals, which can be leveraged for adjusting behavior without re-planning from scratch. While explicit studies on preference-aware recovery in LLM-powered agents remain limited, adjacent work on mitigating tool overuse under diminishing returns (Qian et al., 2025a) and structured reflection against tool-interaction failures (Su et al., 2025) provides foundational mechanisms upon which personalized recovery strategies can be built.
@@ 6.2.2. Outcome Realization
Even when multiple execution outcomes are functionally valid, user satisfaction depends on how results are presented in accordance with individual preferences. In recommendation-oriented agents, personalized re-ranking is a core mechanism for transforming candidate results into ordered outputs that reflect user-specific utility (Pei et al., 2019). Methods such as PEAR and multi-level interaction reranking (MIR) explicitly model cross-item dependencies together with user preference features, yielding more faithful final rankings and improved utility metrics (Li et al., 2022; Xi et al., 2022). Beyond ranking, personalization also governs which execution outcomes are retained or discarded. WARPP (Mazzolenis and Zhang, 2025) emphasizes runtime personalization in workflow execution by pruning or adjusting conditional branches based on user attributes, reducing irrelevant or erroneous outcomes without re-planning the entire process. In addition to selecting among existing results, outcome realization may involve proactive augmentation. ETAPP (Hao et al., 2025) treats proactivity as a personalization signal, evaluating whether agents can propose additional actions or suggestions to better satisfy user needs.
6.3. Discussion
Role of Personalized Action Execution
Action execution constitutes a core stage that bridges high-level reasoning with tangible outcomes in LLM-based agents (Qu et al., 2025; Chowa et al., 2026). By determining how abstract decisions are instantiated, execution-level choices directly influence whether personalization is effectively realized rather than remaining at the level of internal reasoning, which substantially shapes user experience, trust, and perceived utility. Consequently, personalized action execution functions as the final bottleneck of personalization in agent systems, translating inferred user intent into observable behavior and closing the loop between internal personalization mechanisms and external outcomes.
Challenges
Designing personalized action execution mechanisms remains challenging due to several factors. First, execution contexts are highly heterogeneous across tools, environments, and domains, complicating the design of general mechanisms that consistently respect individual preferences. Second, personalization signals at execution time are often subtle, requiring abstract preferences to be translated into concrete parameters rather than explicit instructions. Third, execution feedback may expose ambiguity or brittleness in how user preferences were interpreted upstream, necessitating careful handling and, in some cases, coordination with planning or preference modeling components to avoid cascading errors.
@@ 6.3. Discussion
Research on personalized action execution remains relatively limited, leaving several promising directions open. At the execution primitive level, developing personalization-aware tools that emphasize non-functional characteristics could enable differential behavior beyond task correctness. At the control level, learning preference-conditioned execution policies that generalize across tools and environments may allow agents to reuse personalization signals more effectively. Additionally, agents should better handle vague or missing arguments, localize execution failures, and aggregate outcomes in ways that reflect user priorities rather than default heuristics. Finally, there is a growing need for execution-level evaluation protocols that move beyond task success to assess preference adherence, consistency, and user satisfaction, in line with recent calls for more holistic evaluation of agent capabilities.
@@ 7. Evaluation
Evaluating personalized agents presents a unique set of challenges that extend beyond the traditional assessment of AI systems, as the target of optimization is not merely task correctness but long-horizon user-specific satisfaction. These challenges largely stem from the implicit, context-dependent, and evolving nature of human preferences and desires, which are often under-specified, situation-sensitive, and updated through interaction. Consequently, evaluation should prioritize how well an agent aligns with an individual user’s expectations and experience, rather than only how well it completes a task in an objective sense. This motivates the need for a comprehensive evaluation mechanism that characterizes personalization from general principles down to “in-one’s-shoes” assessments that reflect each user’s subjective utility. In this section, we provide a systematic overview of evaluation for personalized LLM-powered agents, with Figure 5 summarizing how evaluation goals, assessment paradigms, and benchmark families are interconnected within a unified framework.
@@ 7.1. Evaluation Goals and Metrics
To rigorously evaluate personalized agents, we must move beyond conventional task-centric metrics and adopt an evaluation framework that reflects the inherently user-contingent nature of personalization. Unlike standard AI systems, where performance is primarily measured by objective task success, personalized agents are optimized for subjective utility: their behavior should align with an individual user’s preferences, remain consistent across contexts and over time, and adapt as preferences are revealed or revised through interaction, while operating within safety and privacy constraints. Motivated by these considerations, we organize evaluation metrics for personalized agents into five complementary dimensions: Effectiveness, Adaptivity, Generalization, Robustness, and Risk, as illustrated in Table 1. Together, these dimensions provide a structured and holistic view of personalization quality, enabling systematic comparison across diverse tasks, preference sources, and evaluation protocols.
@@
Discovery Accuracy (Li et al., 2025g; Jiang et al., 2024a) Measures whether the agent can correctly discover user preferences or intents from the interaction.
@@
Preference Alignment (Zhao et al., 2025a; Wu et al., 2025e) Measures the alignment between the agent’s output and user preferences or stated constraints.
@@
Preference Lift (Li et al., 2025g; Hao et al., 2025; Wu et al., 2024c) Compares personalized outputs against a non-personalized baseline on the same prompt, reporting the lift in preference score.
@@
Emotional Empathy (Wu et al., 2025e) Measures the degree to which the response shows understanding and compassion for the user’s emotional state.
@@
Adaptivity Interaction Efficiency (Li et al., 2025g) Measures the efficiency of adapting, such as number of turns, clarification questions, or edits needed to satisfy user expectations.
Proactivity (Hao et al., 2025) Measures the ability of the agent to proactively identifying user needs and offering extra suggestions or actions to enhance user satisfaction.
Generalization Out-of-Domain Performance (Jiang et al., 2025b) Measures the ability of the agent to generalize user preference to other scenarios or tasks without explicit instruction.
Robustness Accuracy under Ambiguity (Huang et al., 2025d; Pakhomov et al., 2025) Measures the accuracy of the agent’s action when some information is missing or ambiguous.
@@
Safety Rate (Wu et al., 2025b) Measures whether the agent can make safe decisions based on different user profiles.
Risk Risk Sensitivity (Wu et al., 2025e) Measures whether the agent can recognize and respond appropriately to potential risks in the user’s context.
Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) Measures the percentage of conversations where the agent makes disclosure in the tasks related to users’ privacy information.
@@
Effectiveness is the primary criterion for personalization quality and captures whether an agent can deliver user-contingent utility rather than generic helpfulness. It includes (i) preference discovery, where the agent must infer what the user wants from partial interaction traces or contextual signals (Discovery Accuracy (Jiang et al., 2024a)), (ii) preference usage, where agents should correctly incorporates user-specific information into generation (Knowledge Integration Score (Au et al., 2025; Wang et al., 2023a)), and (iii) preference satisfaction, where the final response should comply with explicit constraints and implicit preferences (Preference Alignment (Zhao et al., 2025a; Wu et al., 2025e), Preference-Aware Planning Accuracy (Xu et al., 2025b, c)). Beyond absolute alignment, Preference Lift (Li et al., 2025g; Hao et al., 2025; Wu et al., 2024c) quantifies gains over a non-personalized baseline under the same prompt. In multi-turn settings, Consistency Score (Mukhopadhyay et al., 2025) evaluates stability of persona and values over time, while Emotional Empathy (Wu et al., 2025e) captures affect-appropriate supportive responses.
@@
Adaptivity evaluates whether a personalized agent can update its behavior as new evidence about preferences arrives while preserving continuity and minimizing user burden. Adaptation Success Rate (Shang et al., 2025; Jiang et al., 2025b) measures whether the agent improves personalization after receiving explicit feedback, new constraints, or additional user context. Selective Forgetting (Hu et al., 2025) that measures the ability to overwrite or retract outdated memory entries when faced with contradictory evidence is equally important due to the evolution of user preferences. Interaction Efficiency (Li et al., 2025g) captures the cost of adaptation, such as the number of turns, clarifications, or edits needed before the user accepts the result, which directly impacts practical usability. Proactivity (Hao et al., 2025) complements these metrics by measuring whether the agent can anticipate needs and offer helpful next steps at the right time without introducing irrelevant suggestions or unnecessary interruptions.
@@
Generalization captures whether personalization transfers beyond the narrow contexts where it was learned. Out-of-Domain Performance (Jiang et al., 2025b) measures whether an agent can apply a user’s preferences in specific scenarios to new domains, tasks, or situations without requiring the user to restate them. This is particularly important for agents that aim to support diverse daily activities, where preferences learned in one domain should carry over to other domains in a principled way.
@@
Personalized agents operate under inherently noisy and incomplete preference evidence, so a useful evaluation must measure reliability under distributional stress rather than only average-case alignment. Two robustness pressures are especially consequential. Firstly, user preferences are frequently underspecified, partially observed, or expressed indirectly. Therefore, Accuracy under Ambiguity (Huang et al., 2025d; Pakhomov et al., 2025) evaluates whether the agent can make reasonable inferences, ask targeted clarification questions, and avoid overconfident assumptions when key signals are missing. Secondly, user inputs can be wrong or based on misconceptions, either unintentionally or due to stale information. Misinformed Condition Performance (Huang et al., 2025a) captures whether the agent can identify incorrect premises, provide corrective guidance, and still preserve a personalized interaction style while avoiding mistaken beliefs.
@@
Risk metrics quantify safety, privacy, and other downside constraints that are inseparable from personalization. Key metrics include: Safety Rate (Wu et al., 2025b) measures whether the agent remains safe across heterogeneous user profiles, including cases where preferences might conflict with policies or where the user requests risky actions. Risk Sensitivity (Wu et al., 2025e) evaluates whether the agent recognizes contextual hazards and calibrates its advice, for example by providing warnings, suggesting safer alternatives, or encouraging professional help when appropriate. Because personalization relies on user-specific information, Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) measures whether the agent discloses private data in disallowed ways, especially under adversarial or indirect prompts. Conversely, Over-Secrecy Rate (Mukhopadhyay et al., 2025) measures the utility loss from excessive caution, where the agent refuses or withholds benign, helpful information despite the user’s legitimate request. Together, these metrics target the core trade-off in personalized agents that aims to maximize user utility subject to risk constraints.
@@ 7.2. Assessment Paradigms
The above metrics are instantiated through different assessment paradigms depending on whether the target criterion is objectively verifiable or user-contingent. We group mainstream paradigms into four families. (1) Automatic scoring applies when ground-truth labels or reference answers exist, for example accuracy, precision/recall/F1 scores for classification, and overlap-based metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) for generation. (2) Rule-based constraint checking evaluates compliance with verifiable rubrics or hard constraints, such as whether required attributes are included, forbidden attributes are avoided, or preference constraints are satisfied in structured outputs. This protocol is particularly useful for delegation-style tasks where constraints can be deterministically validated. (3) Learned LLM-based evaluators (LLM-E) fine-tune evaluation models to provide more stable, fine-grained diagnostic feedback for specific preference dimensions, trading off generality for higher consistency in the target domain (Wang et al., 2024a). (4) LLM-as-a-judge (LLM-J) is commonly used for user-contingent criteria such as preference alignment, tone, and satisfaction, where the judge model is conditioned on a user profile or preference description and asked to rate or rank candidate outputs (Zheng et al., 2023). To improve reliability, LLM-J evaluations often adopt pairwise comparisons, calibrated rubrics, and multi-judge aggregation.
@@ 7.3.1. Interactive Alignment Benchmarks
A core challenge in interactive alignment is preference discovery where the agent must recognize, infer, and iteratively refine user intents and preferences from heterogeneous signals, ranging from explicit instructions to implicit feedback and contextual cues (Tsaknakis et al., 2025). In this category, IndieValueCatalog (Jiang et al., 2024a) curates value-expressing statements from diverse individuals and evaluates whether models can infer individualized value judgments, highlighting limitations in capturing fine-grained personal values. PersonaBench (Tan et al., 2025a) constructs synthetic yet realistic personal profiles and associated artifacts to test whether agents can extract and reason over user-specific information. PrefDisco (Li et al., 2025g) proposes a meta-evaluation framework that emphasizes interactive elicitation efficiency, measuring whether an agent can proactively ask informative questions for a user-aligned solution with minimal interaction turns.
@@
Complementary to the above, component-probing benchmarks provide controlled testbeds that isolate specific personalization capabilities, enabling diagnostic analysis of when and how an agent leverages user information. PersonaFeedback (Tao et al., 2025) pairs predefined personas with queries and asks the model to select persona-consistent responses, offering a direct probe of preference-conditioned alignment. A major sub-line focuses on personalized memory, including LongMemEval (Wu et al., ), LoCoMo (Maharana et al., 2024), PerLTQA (Du et al., 2024), MemoryAgentBench (Hu et al., 2025), ConvoMem (Pakhomov et al., 2025), and PAL-Bench (Huang et al., 2025f), which test whether agents can organize, retrieve, and apply user-specific memories under long conversations or incremental preference revelation. For personalized tool use, ETAPP (Hao et al., 2025) and PTBench (Huang et al., 2025d) evaluate whether agents tailor tool-invocation strategies to user needs, including proactivity and handling missing or ambiguous information. Some benchmarks explicitly target adaptivity over time; for example, PersonaMem and PersonaMem-v2 (Jiang et al., 2025a, b) assess whether an agent can track evolving traits and preferences, revise outdated beliefs, and generalize them to new scenarios. Finally, several benchmarks diagnose user-specific risk, such as PSG-Agent (Wu et al., 2025b) and PENGUIN (Wu et al., 2025e) that evaluate profile-contingent safety considerations for LLM-based agents, while PrivacyBench (Mukhopadhyay et al., 2025) and AgentDam (Zharmagambetov et al., 2025) test privacy compliance for agents when handling sensitive user information.

### v2

@@ 6.2. Post-action Stage
The post-action stage closes the loop after an action is executed. Given execution result, the agent evaluates whether the outcome satisfies user-conditioned objectives, applies corrective adjustments when mismatches arise (execution-time correction), and produces final outputs that conform to user preferences and quality criteria (outcome realization).
@@ 6.2.1. Execution-Time Correction
A failed execution commonly produces feedback signals, which can be leveraged for adjusting behavior without re-planning from scratch. Explicit studies on preference-aware recovery in PLAs remain limited. Nevertheless, adjacent work suggests useful mechanisms for this stage. For example, research on mitigating tool overuse under diminishing returns (Qian et al., 2025a) and structured reflection over tool-interaction failures (Su et al., 2025) provides building blocks for future personalized recovery strategies.
@@ 6.2.2. Outcome Realization
Even when multiple execution outcomes are functionally valid, user satisfaction depends on how final results are presented in accordance with individual preferences. One common mechanism is personalized re-ranking, which transforms candidate results into outputs that better reflect user-specific utility (Pei et al., 2019). Methods such as PEAR and MIR explicitly model user preference features together with cross-item dependencies, producing more faithful final rankings (Li et al., 2022; Xi et al., 2022). Personalization may also govern which outcomes are retained or discarded: WARPP (Mazzolenis and Zhang, 2025) adjusts workflow branches at runtime based on user attributes, reducing irrelevant or erroneous outcomes without re-planning the full process. Finally, outcome realization can extend beyond selecting among existing results to proactively augmenting them. ETAPP (Hao et al., 2025) treats proactivity as a personalization signal, evaluating whether agents can propose additional actions or suggestions to better satisfy user needs.
Role and Challenges.
Action execution is the stage at which personalized decisions are translated into concrete operations and external outcomes in PLAs (Qu et al., 2025; Chowa et al., 2026). As such, it bridges high-level reasoning with observable behavior, determining whether personalization is actually realized rather than remaining at the level of internal reasoning. This stage is especially important because user experience depends not only on what the agent decides, but also on how those decisions are instantiated and delivered. At the same time, personalized action execution remains challenging for several reasons. Execution contexts are highly heterogeneous across tools, environments, and domains, making it difficult to design general mechanisms that consistently respect individual preferences. Moreover, execution-time personalization signals are often subtle, requiring abstract preferences to be translated into concrete parameters rather than explicit instructions. Finally, execution feedback may reveal ambiguity or brittleness in how user preferences were interpreted upstream, requiring careful handling and, in some cases, coordination with planning or preference-modeling components to avoid cascading errors.
@@
Research on personalized action execution remains relatively limited, leaving several promising directions open. One is the development of personalization-aware execution primitives and tools that account not only for task correctness but also for non-functional user preferences. Another is learning preference-conditioned execution policies that generalize across tools and environments, enabling agents to reuse personalization signals more effectively. It is also important to improve how agents handle vague or missing arguments, localize execution failures, and aggregate outcomes in ways that reflect user priorities rather than default heuristics. More broadly, this stage would benefit from execution-level evaluation protocols that move beyond task success to assess preference adherence, consistency, and user satisfaction.
@@ 7. Evaluation
Evaluating personalized agents requires going beyond conventional task-centric assessment, because the objective is not only task correctness but also long-horizon, user-specific utility. This challenge stems from the fact that human preferences are often implicit, context-dependent, and evolving through interaction, making personalization difficult to measure with a single objective criterion. As a result, evaluation should capture not only whether an agent completes a task, but also whether it aligns with an individual user’s preferences, expectations, and experience over time. In this section, we provide a systematic overview of evaluation for personalized LLM-powered agents. As summarized in Figure 5, we organize the discussion into three layers: evaluation goals and metric dimensions, assessment paradigms, and representative benchmark families.
@@ 7.1. Evaluation Goals and Metrics
A rigorous evaluation framework for personalized agents must reflect the multi-faceted nature of personalization quality. Beyond objective task success, an agent should align with individual preferences, remain coherent across contexts and over time, adapt when preferences are revealed or revised, and operate within safety and privacy constraints. Motivated by these requirements, we organize evaluation metrics for personalized agents into five complementary dimensions: Effectiveness, Adaptivity, Generalization, Robustness, and Risk, as illustrated in Table 2. These metric dimensions do not apply uniformly across all capabilities, but provide a structured basis for comparing personalized agents across tasks, preference sources, and assessment paradigms.
Table 2. Evaluation metrics for personalized LLM-powered agents.
@@
Discovery Accuracy (Li et al., 2025h; Jiang et al., 2024a) Measures whether the agent can correctly discover user preferences or intents from the interaction.
@@
Preference Alignment (Zhao et al., 2025a; Wu et al., 2025f) Measures the alignment between the agent’s output and user preferences or stated constraints.
@@
Preference Lift (Li et al., 2025h; Hao et al., 2025; Wu et al., 2024c) Compares personalized outputs against a non-personalized baseline on the same prompt, reporting the lift in preference score.
@@
Emotional Empathy (Wu et al., 2025f) Measures the degree to which the response shows understanding and compassion for the user’s emotional state.
@@
Adaptivity Interaction Efficiency (Li et al., 2025h) Measures the efficiency of adapting, such as number of turns, clarification questions, or edits needed to satisfy user expectations.
Proactivity (Hao et al., 2025) Measures the ability of the agent to proactively identify user needs and offering extra suggestions or actions to enhance user satisfaction.
Generalization Out-of-Domain Performance (Jiang et al., 2025b) Measures the ability of the agent to generalize user preferences to other scenarios or tasks without explicit instruction.
Robustness Accuracy under Ambiguity (Huang et al., 2025c; Pakhomov et al., 2025) Measures the accuracy of the agent’s action when some information is missing or ambiguous.
@@
Safety Rate (Wu et al., 2025c) Measures whether the agent can make safe decisions based on different user profiles.
Risk Risk Sensitivity (Wu et al., 2025f) Measures whether the agent can recognize and respond appropriately to potential risks in the user’s context.
Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) Measures the percentage of conversations in which the agent discloses private information inappropriately in tasks related to users’ privacy information.
@@
Effectiveness captures whether a personalized agent can produce user-contingent utility rather than merely generic helpfulness. It covers at least three aspects: whether the agent can infer what the user wants from partial interaction traces or contextual signals; whether it can correctly incorporate user-specific information into planning and generation; and whether the final response or action actually satisfies explicit constraints and implicit preferences. Metrics such as Discovery Accuracy (Jiang et al., 2024a), Knowledge Integration Score (Au et al., 2025; Wang et al., 2023a), Preference Alignment (Zhao et al., 2025a; Wu et al., 2025f), and Preference-Aware Planning Accuracy (Xu et al., 2025b, c) target these aspects from different angles. Beyond absolute alignment, Preference Lift (Li et al., 2025h; Hao et al., 2025; Wu et al., 2024c) measures gains over a non-personalized baseline under the same prompt. In multi-turn settings, Consistency Score (Mukhopadhyay et al., 2025) evaluates stability of persona and values over time, while Emotional Empathy (Wu et al., 2025f) captures whether responses remain affect-appropriate and supportive.
@@
Adaptivity evaluates whether a personalized agent can update its behavior as new evidence about user preferences becomes available, while preserving continuity and minimizing user burden. Adaptation Success Rate (Shang et al., 2025; Jiang et al., 2025b) measures whether the agent improves personalization after receiving explicit feedback, additional constraints, or new user context. Because preferences may evolve or even contradict earlier evidence, Selective Forgetting (Hu et al., 2025) is equally important for assessing whether outdated memory can be revised, overwritten, or removed appropriately. Interaction Efficiency (Li et al., 2025h) captures the cost of adaptation, such as the number of turns, clarifications, or edits required before the user accepts the result. Proactivity (Hao et al., 2025) complements these metrics by measuring whether the agent can anticipate needs and offer helpful next steps at the right time without creating unnecessary interruption or irrelevant suggestions.
@@
Generalization captures whether personalization transfers beyond the contexts in which it was originally learned. In practice, a useful personalized agent should not require the user to restate preferences for every new domain, task, or situation. Out-of-Domain Performance (Jiang et al., 2025b) therefore measures whether preferences acquired in one setting can be applied appropriately in others. This dimension is especially important for general-purpose agents that are expected to support diverse daily activities, where personalization must remain reusable rather than narrowly tied to a single task context.
@@
Personalized agents operate under noisy, incomplete, and sometimes misleading preference evidence, so evaluation must measure reliability under stress rather than only average-case alignment. Two robustness pressures are especially important. First, user preferences are often underspecified, partially observed, or expressed indirectly. Accuracy under Ambiguity (Huang et al., 2025c; Pakhomov et al., 2025) therefore evaluates whether the agent can make reasonable inferences, ask targeted clarification questions, and avoid overconfident assumptions when key signals are missing. Second, user inputs may be incorrect or based on misconceptions. Misinformed Condition Performance (Huang et al., 2025a) measures whether the agent can identify flawed premises, provide corrective guidance, and still preserve a personalized interaction style.
@@
Risk metrics quantify safety, privacy, and other downside constraints that are inseparable from personalization. Safety Rate (Wu et al., 2025c) measures whether the agent remains safe across heterogeneous user profiles, including cases where user preferences may conflict with policies or where requests involve risky actions. Risk Sensitivity (Wu et al., 2025f) evaluates whether the agent can recognize contextual hazards and calibrate its advice appropriately, for example, by providing warnings, suggesting safer alternatives, or encouraging professional help when needed. Because personalization relies on user-specific information, Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) measures whether the agent discloses private data in disallowed ways, especially under adversarial or indirect prompts. Conversely, Over-Secrecy Rate (Mukhopadhyay et al., 2025) captures utility loss from excessive caution, where the agent withholds benign but helpful information despite a legitimate request. These metrics reflect the central challenge of maximizing user utility while respecting risk constraints.
@@ 7.2. Assessment Paradigms
The metric dimensions above can be operationalized through different assessment paradigms, depending on whether the target criterion is objectively verifiable or inherently user-contingent. We group mainstream paradigms into four families. (1) Automatic scoring applies when ground-truth labels or reference answers are available. Typical examples include accuracy and precision/recall/F1 for classification, as well as overlap-based metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) for generation. (2) Rule-based constraint checking evaluates whether outputs satisfy explicit and verifiable requirements, such as including required attributes, avoiding forbidden ones, or respecting structured preference constraints. This paradigm is especially useful for delegation-style tasks in which compliance can be deterministically validated. (3) Learned LLM-based evaluators (LLM-E) train dedicated evaluation models to assess specific preference dimensions, often providing more stable and fine-grained diagnostic feedback at the cost of reduced generality (Wang et al., 2024a). (4) LLM-as-a-judge (LLM-J) uses a general-purpose LLM as the evaluator for user-contingent criteria such as preference alignment, tone, and satisfaction. In this setting, the judge is typically conditioned on a user profile or preference description and asked to score or rank candidate outputs (Zheng et al., 2023). To improve reliability, LLM-J protocols often adopt pairwise comparison, calibrated rubrics, and multi-judge aggregation.
@@ 7.3.1. Interactive Alignment Benchmarks
A core challenge in interactive alignment is preference discovery, where the agent must recognize, infer, and iteratively refine user intents and preferences from heterogeneous signals, ranging from explicit instructions to implicit feedback and contextual cues (Tsaknakis et al., 2025). In this category, IndieValueCatalog (Jiang et al., 2024a) curates value-expressing statements from diverse individuals and evaluates whether models can infer individualized value judgments, highlighting limitations in capturing fine-grained personal values. PersonaBench (Tan et al., 2025a) constructs synthetic yet realistic personal profiles and associated artifacts to test whether agents can extract and reason over user-specific information. PrefDisco (Li et al., 2025h) proposes a meta-evaluation framework that emphasizes interactive elicitation efficiency, measuring whether an agent can proactively ask informative questions for a user-aligned solution with minimal interaction turns.
@@
Complementary to the above, component-probing benchmarks provide controlled testbeds that isolate specific personalization capabilities, enabling diagnostic analysis of when and how an agent leverages user information. PersonaFeedback (Tao et al., 2025) pairs predefined personas with queries and asks the model to select persona-consistent responses, offering a direct probe of preference-conditioned alignment. A major sub-line focuses on personalized memory, including LongMemEval (Wu et al., 2025a), LoCoMo (Maharana et al., 2024), PerLTQA (Du et al., 2024), MemoryAgentBench (Hu et al., 2025), ConvoMem (Pakhomov et al., 2025), and PAL-Bench (Huang et al., 2025d), which test whether agents can organize, retrieve, and apply user-specific memories under long conversations or incremental preference revelation. For personalized tool use, ETAPP (Hao et al., 2025) and PTBench (Huang et al., 2025c) evaluate whether agents tailor tool-invocation strategies to user needs, including proactivity and handling missing or ambiguous information. Some benchmarks explicitly target adaptivity over time; for example, PersonaMem (Jiang et al., 2025a, b) assess whether an agent can track evolving traits and preferences, revise outdated beliefs, and generalize them to new scenarios. Finally, several benchmarks diagnose user-specific risk, such as PSG-Agent (Wu et al., 2025c) and PENGUIN (Wu et al., 2025f) that evaluate profile-contingent safety considerations for LLM-based agents, while PrivacyBench (Mukhopadhyay et al., 2025) and AgentDam (Zharmagambetov et al., 2025) test privacy compliance for agents when handling sensitive user information.